OpenStudy (anonymous):

A printer numbers the pages of a book starting with 1 and uses 3189 digits in all. How many pages does the book have ?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

how many pages does the book have

OpenStudy (anonymous):

1074

OpenStudy (anonymous):

I don't need the answer , i need explanation

OpenStudy (anonymous):

No. of digits in 1-digit page nos. = 1x9 = 9. No. of digits in 2-digit page nos. = 2 x 90 = 180. No. of digits in 3-digit page nos. = 3 x 900 = 2700. No. of digits in 4-digit page nos. = 3189 - (9 + 180 + 2700) = 3189 - 2889 = 300. Therefore No. of pages with 4-digit page nos. = (300/4) = 75. Hence, total number of pages = (999 + 75) = 1074.

OpenStudy (anonymous):

I know that is from a website , but i didn't get it though

OpenStudy (anonymous):

really

OpenStudy (anonymous):

then i dont know

OpenStudy (anonymous):

sry

OpenStudy (anonymous):

thank you though for helping